OpenStudy (anonymous):

a farmer feeds his cows 200 pounds of feed each day and has 700 pounds of feed in his barn. Another farmer feeds his cows 350 pounds of feed each day and has 1000 pounds of feed in his barn. in how many days will the two farmers have the same amount of feed left? how would your answer change if both farmers got an additional 1000 pounds of feed? plz help don't know how to set up the equations...

OpenStudy (krazykricket):

Set the problem up like this: 700−200x=1000−350x⟹150x=300⟹x=2 In two days, both farmers will have four hundred pounds of feed.

OpenStudy (krazykricket):

The adjusted problem for 1000 lb more at each farm is similar: 1700−200x=2000−350x Amazingly (or maybe not.....) the answer is still two days, although they will each have 1400 lb instead of 400 lb.

OpenStudy (anonymous):

thx ^^^ @KrazyKricket

OpenStudy (krazykricket):

Your Welcome